3. Two liters of the known carcinogen, carbon tetrachloride (CCI4), are spilled into a room (100-m x 200-m x 50-m). Determine the equilibrium concentration of CCl4 in the air at 220C. For CC PO -0.130 atm, p 1.59 g/mL, Henrys Law Coefficient 3.04 x 10 2atm.m3/molSolve without using the ideal gas law
Answer #1

From Henry's law

P =KH * C

P = partial pressure = 0.130 atm

KH = Henry's law constant = 3.04*10-2 atm-m3 /mol

Hence, concentration C = P/KH = 0.13 /(3.04*10-2) = 4.2763 mol/m3

Volume of CCl4 = 2 L = 2*10-3 m3

Moles of CCL4 = 4.2763*2*10-3 = 8.552*10-3 moles initially

Now, volume of room = 100*200*50 = 106 m3

Hence, equilibrium concentration in the room = 8.552*10-3 /(106) = 8.552*10-9 mol/m3